Assume that a firm’s beginning capital stock is $500,000 and capital depreciates at a rate of five percent per annum. Now assume that the firm also spends the following on gross investment (starting from the beginning year): $10,000, $30,000, and $30,000. Then the firm’s capital stock at the beginning of the third year is different from the beginning capital stock by

(1) $70,000.00.

(2) -$25,000.00.

(3) -$3,787.50.

(4) $45,000.00
Answer #1

The capital at the beginning of year 1 = $500,000.

Capital at the end of year 1 = Initiatial capital stock -Depreciation + Investment

Capital at the end of year 1= 500000 -500000*0.05 + 10000 = $485000

Similarly, Capital at the end of year 2 will be = 485000 -485000*0.05 + 30000= $490750

similarly, capital at the end of the year 3 will be = 490750 - 490750*0.05 + 30000 = $496212.5

so the difference between capital at the beginning of the first year and the capital at the end of the third year will be = 496212.5 - 500,000. = -$3,787.50.